[Brahms-l] Shift report 20040305 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 05 2004 - 00:19:09 EST
------------------------------------------------------------
Date: Fri Mar  5 00:19:09 2004, Shift: 16:00-24:00
Supervisor: hiro, with: 
Log book#15 pages 75 - 75

Shift Summary:
------------------------------------------------------------
1.  Machine development until about 20:00
    beam after that.
2.  10759, 10760
    10761 speicial run with H1 out of trigger 2
3.  Disc level of TRFS top 2 changed to 120mV from 50mV during the short access.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 5 00:19:26 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 05 2004 - 00:19:57 EST